OpenStudy (anonymous):

In the triangle below, A = 35°, B = 50°, and a = 12 feet. Find the other values.

OpenStudy (slaaibak):

What triangle?

OpenStudy (anonymous):

simaler figure?

OpenStudy (slaaibak):

C = 95 degrees

OpenStudy (anonymous):

law of sines for this one use 16 y^3 (y^4+2 y^2+5) \[\frac{b}{\sin(50)}=\frac{12}{\sin(35)}\] to solve for b

OpenStudy (anonymous):

so \[b=\frac{12\sin(50)}{\sin(35)}\] then a calculator

OpenStudy (king):

third angle is 95

OpenStudy (anonymous):

hey thanks i have that same question!!!

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

i get \[b=16\] rounded

OpenStudy (anonymous):

thank you,thank you!

OpenStudy (king):

ya me 2!16

OpenStudy (anonymous):

that was wrong. repeat with \[c=\frac{12\sin(95)}{\sin(35)}\]

OpenStudy (anonymous):

looks like [c=20.9\]rounded. makes sense because the longest side is opposed the largest angle

OpenStudy (anonymous):

The other values are b = 16.03 feet, c = 20.84 feet and c = 95.

OpenStudy (anonymous):

dang your soo smart....